solve the system of equation by graphing.
x+y=7
y-x=7
The solution set of the system of equation is

Answer by Alan3354(69443)

solve the system of equation by graphing.
x+y=7
y-x=7
---------
They're 2 straight lines.
For x+y = 7
y = -x + 7
Pick 2 values for x and find y
Plot the 2 points.
Draw a line thru them.
----
Do the same for the other eqn.
The intersection is the solution of the system.